Wind blows along S-E direction with a speed of 5m/s.A plstic bag is thrown with a velocity of 5(i+2j+k) m/s from a building 10m high then the position vector of the point where the bag lends off on the earth will be(x,y and z axes respectively point along east,north and vertically up directions, the point of projection being origin)_____________

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sindhuja

velocity of wind =5/√2 (i-j)

velocity of plastic bag =5(i+2j+k)

resultant velocity of plastic bag =5(i+2j+k) +5/√2 (i-j)

                                                           =(5+5/√2)i +(10-5/√2)j +5k

time take to reach the ground

10=5t +1/2 *10* t2

t2 +t -1=0

t=1 sec

in this time distance travel in x direction =(5+5/√2) *1 =5+5/√2

in this time distance travel in y direction =(10-5/√2) *1 =10-5/√2

so position vector  =(5+5/√2)i +(10-5/√2)j +0k


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin